List all the solutions of the equation:
the absolute value of 2y-3 =9

Answer by lwsshak3(11628)

List all the solutions of the equation:
the absolute value of 2y-3 =9
**
|2y-3|=9
for (2y-3)>0
2y-3=9
2y=12
y=6
..
for (2y-3)<0
-|2y-3|=9
-2y+3=9
-2y=6
y=-3